In a certain city you want to estimate the percentage of households that use oil as fuel for heating. What size should the sample be? if you want to have a confidence of 98% with a margin of error of 3 percentage points.

a) 1503

b) 1309

c) 1172

d) 1068
Answer #1

p(hat)= 0.5 and q(hat)= 1-0.5= 0.5

n= p(hat)*q(hat)*Zc^2,/E^2

n= 0.5*0.5*(2.33)^2/(0.03)^2

n= 1503 OPTION A